# AD-GS: Object-Aware B-Spline Gaussian Splatting for Self-Supervised Autonomous Driving

### v1

@@ 3.2 Learnable B-Spline Motion Curve
where $t_{0},t_{1},...,t_{n+1}$ are the knots arranged in non-decreasing order. In practice, we define $t$ as the timestamp, with $t_{k-1}$ and $t_{n+1}$ representing the minimum and maximum timestamps of the input sequence, respectively. The basis function $B_{i,k}$ only has a non-zero value in $[t_{i},t_{i+k}]$. Furthermore, all control points $\mathbf{p}_{i}$ are treated as learnable parameters, and we utilize uniform B-spline curves for motion modeling. However, the de Boor-Cox recursion formula for the basis function $B_{i,k}$ is computationally inefficient. To address this, we employ the matrix formulation of uniform B-spline curves, which offers better computational complexity [24].
@@ 3.3 Object-Aware Splatting with Temporal Mask
where $\mathbb{I}$ denotes the indicator function. The rendered object mask is supervised via the pseudo ground truth through the binary cross-entropy loss $\mathcal{L}_{obj}=BCE(\hat{\mathcal{M}}_{obj},\mathcal{M}_{obj})$. During training, the loss $\mathcal{L}_{obj}$ ensures that the two Gaussian subsets are maintained in their respective areas by adjusting their opacity. More details can be found in Section A of our supplementary.
Bidirectional Temporal Visibility Mask. The moving objects might not be visible in all frames, necessitating the use of a temporal mask to account for their sudden appearance and disappearance. Therefore, we introduce the bidirectional temporal visibility mask applied to the opacity $\sigma$ of an object Gaussian $G\in\boldsymbol{\Omega}_{obj}$, defined by the formula:
@@
where $\mu_{t}$ is the fixed acquisition timestamp of the LiDAR points, $s_{0},s_{1}$ are the learnable scales of the bidirectional visibility mask. Unlike previous approaches [2], which treat $\mu_{t}$ as a learnable parameter, we argue that the acquisition timestamp itself serves as an effective prompt for determining the visible moment of an object. To prevent the temporal mask from becoming excessively narrow, we design an expanding regularization loss as follows:
@@ 3.4 Self-Supervision with Regularization
where $\boldsymbol{\mu}^{\prime}$ is the position of each Gaussian at the target moment. We then project the $\hat{X}_{r}$ onto the target image, and supervise it using the corresponding pseudo flow points $X_{t}$ through $\mathcal{L}_{f}=||\phi_{t}(\hat{X}_{r})-X_{t}||_{1}$, where $\phi_{t}$ is the projection function. To enhance efficiency and reduce noise, we only supervise the object part of the flow according to the simplified 2D pseudo segmentation results obtained from SAM [25, 11].
Inverse Depth Supervision. In prior work, the depth map is rendered through the $\alpha$-blending of the distances between the camera and Gaussian centers, which is then inverted to supervise against the monocular pseudo ground truth. However, when no Gaussian exists along a pixel ray, the inversion operation can become computationally unstable. To address this, we propose directly rendering the expectation map of inverse depth by using the following formula:
@@ 4.3 Ablation Study
Ablation of Modules. To demonstrate the effectiveness of each module in AD-GS for object modeling, we perform an ablation study by progressively incorporating the learnable B-spline curves and the bidirectional temporal visibility masks, corresponding to “B-spline” and “t-mask”, respectively. For the “sin&cos” setting, the B-spline quaternion curves are replaced by the trigonometric functions. The results are presented in Table 5 and Figure 8. Only using B-splines can result in over-fitting due to the noisy self-supervision, and trigonometric functions reduce the noise through their global optimization properties. The last row of Figure 8 illustrates the scenarios where a car suddenly appears and disappears. Although the trigonometric functions successfully capture the general motion under the noisy self-supervision, the representation is inaccurate due to the incorrect optimization during the nearby training frames where the car is invisible. The B-spline curves provide local motion fitting for more accurate representation to reduce the negative effect from the invisible frames, and the bidirectional temporal visibility masks further mitigate this problem. More ablations studies can be found in Section C of our supplementary.
@@ Appendix C Additional Results
Analysis of Motion Fitting. To further demonstrate the effectiveness of trigonometric function and B-spline curve in global and local fitting, we visualize the deformation map mainly following the approach in Grid4D [33]. The results are shown in Figure 9, where similar colors indicate similar deformations. Although trigonometric functions can approximate the general deformation of an object under the noisy self-supervision, their representation tends to be inaccurate due to the omission of per-frame local details. In contrast, B-spline curves offer advantages in capturing local details, allowing for more precise fine-tuning of the representation. The bottom scenario in Figure 9 illustrates a special case where a car suddenly appears and then disappears, remaining visible for only about 17 frames (total about 160 frames). In such cases, trigonometric functions might be influenced by numerous invisible training frames, leading to incorrect gradients during optimization and resulting in severe blurring in motion representation and rendering. However, B-spline curves mitigate this issue by optimizing only the relevant control points, thereby reducing the impact of incorrect gradients from invisible training frames and significantly improving the accuracy of the representation. Therefore, by combining trigonometric functions and B-spline curves for motion fitting, we achieve more accurate motion representations.

### v2

@@ 3.2 Learnable B-Spline Motion Curve
where $t_{0},t_{1},...,t_{n+1}$ are the knots arranged in non-decreasing order. In practice, we define $t$ as the timestamp, with $t_{k-1}$ and $t_{n+1}$ representing the minimum and maximum timestamps of the input sequence, respectively. The basis function $B_{i,k}$ only has a non-zero value in $[t_{i},t_{i+k}]$. Furthermore, all control points $\mathbf{p}_{i}$ are treated as learnable parameters, and we utilize uniform B-spline curves for motion modeling. However, the de Boor-Cox recursion formula for the basis function $B_{i,k}$ is computationally inefficient. Therefore, we use the matrix formulation of uniform B-spline curves for better computational complexity [25].
@@ 3.3 Object-Aware Splatting with Temporal Mask
where $\mathbb{I}$ denotes the indicator function. The rendered object mask is supervised via the pseudo ground truth using the binary cross-entropy loss $\mathcal{L}_{obj}=BCE(\hat{\mathcal{M}}_{obj},\mathcal{M}_{obj})$. During training, $\mathcal{L}_{obj}$ makes the two Gaussian subsets stayed in their respective areas by adjusting their opacities. More details can be found in Section A of our supplementary.
Bidirectional Temporal Visibility Mask. The moving objects might not be visible in all frames, necessitating a temporal mask to reduce the impact of invisible frames captured the already traversed places when the trajectories have not been fully fitted. Therefore, we introduce the bidirectional temporal visibility mask applied to the opacity $\sigma$ of an object Gaussian $G\in\boldsymbol{\Omega}_{obj}$, defined by the formula:
@@
where $\mu_{t}$ is the fixed acquisition timestamp of the LiDAR points, $s_{0},s_{1}$ are the learnable scales of the bidirectional visibility mask. Unlike previous approaches [2], which treat $\mu_{t}$ as a learnable parameter, we argue that the acquisition timestamp itself serves as an effective prompt to determine the visible moment of an object and distinguish the different objects that have passed through the same place. To prevent the temporal mask from becoming excessively narrow, we design an expanding regularization loss as follows:
@@ 3.4 Self-Supervision with Regularization
where $\boldsymbol{\mu}^{\prime}$ is the position of each Gaussian at the target moment. We then project the $\hat{X}_{r}$ onto the target image, and supervise it using the corresponding pseudo flow points $X_{t}$ through $\mathcal{L}_{f}=||\phi_{t}(\hat{X}_{r})-X_{t}||_{1}$, where $\phi_{t}$ is the projection function. To enhance efficiency and reduce noise, we only supervise the object part of the flow with the simplified 2D pseudo segmentation results obtained from SAM [26, 12].
Inverse Depth Supervision. In prior work, the depth map is rendered through the $\alpha$-blending of the distances between the camera and Gaussian centers, which is then inverted to supervise against the monocular pseudo ground truth. However, when no Gaussian exists along a pixel ray, the inversion operation can become computationally unstable. To address this, we directly render the expectation map of inverse depth [10] by using the following formula:
@@ 4.3 Ablation Study
Ablation of Modules. To demonstrate the effectiveness of each module in AD-GS for object modeling, we perform an ablation study by progressively incorporating the learnable B-spline curves and the bidirectional temporal visibility masks, corresponding to “B-spline” and “t-mask”, respectively. For the “sin&cos” setting, the B-spline quaternion curves are replaced by the trigonometric functions. The results are presented in Table 5 and Figure 8. Only using B-splines can result in over-fitting due to the noisy self-supervision, and trigonometric functions reduce the noise through their global optimization properties. The last row of Figure 8 illustrates the scenarios in which a car suddenly appears and disappears. Although the trigonometric functions successfully capture the general motion under the noisy self-supervision, the representation is inaccurate due to the incorrect optimization during the nearby training frames where the car is invisible. The B-spline curves provide local motion fitting for more accurate representation to reduce the negative effect from the invisible frames, and the bidirectional temporal visibility masks further mitigate this problem. More details and ablation studies can be found in Section C of our supplementary.
@@ Appendix C Additional Results
Analysis of Motion Fitting. To further demonstrate the effectiveness of trigonometric function and B-spline curve in global and local fitting, we visualize the deformation map mainly following the approach in Grid4D [34]. The results are shown in Figure 9, where similar colors indicate similar deformations. Although trigonometric functions can approximate the general deformation of an object under the noisy self-supervision, their representation tends to be inaccurate due to the omission of per-frame local details. In contrast, B-spline curves offer advantages in capturing local details, allowing for more precise fine-tuning of the representation. The bottom scenario in Figure 9 illustrates a special case where a car suddenly appears and then disappears, remaining visible for only about 17 frames (total about 160 frames). When the model has not been fully optimized, the car still appears in the invisible frames. However, the invisible frames cannot provide the correct information for the model to fit the trajectory at their timestamps. In such cases, trigonometric functions might be influenced by numerous invisible training frames, leading to incorrect gradients during optimization and resulting in severe blurring in motion representation and rendering. However, B-spline curves mitigate this issue by optimizing only the relevant control points, thereby reducing the impact of incorrect gradients from invisible training frames and significantly improving the accuracy of the representation. Therefore, by combining trigonometric functions and B-spline curves for motion fitting, we achieve more accurate motion representations.